OpenStudy (perl):

ou have a cup of coffee whose temperature is 200 degrees. You cannot drink it until it reaches 150 degrees. You also have cream which is 40 degrees. The room you are in is 80 degrees. Assuming your coffee cools according to Newton cooling and than when you add you the cream to the coffee the new temperature of the mixture is instantly 11/13 times the temperature of the coffee plus 2/13 times 40 (the temperature of the cream). Note that the cream isn't warming up because you leave in the refrigerator. Assuming the cooling constant for the coffee is k = .6 minutes^-11, when should you add the

OpenStudy (perl):

continued...when should you add the creamer to ensure you can drink your coffee the soonest

OpenStudy (perl):

so i started it, newtons law of cooling gives us y' = k ( y-80)

OpenStudy (anonymous):

i am stuck all i know is the temperature of the coffee is \(120e^{-.6t}+80\)

OpenStudy (perl):

right thats what i got too

OpenStudy (perl):

the solution is given here, but i dont understand it http://math.stackexchange.com/questions/363066/newton-cooling-coffee

OpenStudy (anonymous):

although i have no idea what \(\text{minutes}^{-11}\) means

OpenStudy (perl):

maybe thats a typo

OpenStudy (anonymous):

wow that stack exchange is a lot of words

OpenStudy (anonymous):

Just looking at that, too @perl . So, what that answer says is that, using the theory, we want to find a time \(s\) such that, when the cream is poured in, the coffee is immediately drinkable, which is a reasonable reduction based on the problem, as the time per Centigrade/Fahrenheit extends as such \(t\) increases.

OpenStudy (perl):

@LolWolf right, we want to add cream so that the coffee reaches 150 degrees soonest (so its drinkable)

OpenStudy (anonymous):

i think the idea is now to put \[150=\frac{2}{13}\times 40+\frac{11}{13}(80+120e^{-.6t})\] and solve for \(t\)

OpenStudy (anonymous):

Yeah, the simplification, though, comes in the case that it must be \immediately\ drinkable afterwards. So, find a time t, such that, when the cream is added, it becomes immediately below the drinkable-temperature threshold.

OpenStudy (perl):

but where do we take the derivative, it says to minimize?

OpenStudy (anonymous):

The minimization isn't necessary under the assumptions.

OpenStudy (anonymous):

the "minimize" is the part we you say intuitively it cools fastest at the beginning, add the cream at the end

OpenStudy (perl):

oh so we dont need to take the derivative

OpenStudy (anonymous):

I mean, if you want to be rigorous, then sure, you can show this, but I don't believe it's necessary for the answer.

OpenStudy (perl):

I took the derivative of y(T) = 2/13*40 + 11/13(80 + 120 e^(-.6T))

OpenStudy (perl):

but that didnt work

OpenStudy (perl):

ok let me solve your equation

OpenStudy (perl):

after simplifying i get 3/4 = e^(-.6t) ln(3/4)/(-.6) = t

OpenStudy (perl):

so at t = .47947 minutes approximately, we add the cream

OpenStudy (perl):

so the calculus part was getting that equation y = 120 e^(-.6t) + 80

OpenStudy (perl):